Renee’s Boutique, Inc., needs to raise $58.20 million to finance firm expansion. In discussions with its investment bank, Renee’s learns that the bankers recommend a debt issue with an offer price of $1,000 per bond and they will charge an underwriter’s spread of 8.0 percent of the gross price.

Calculate the net proceeds to Renee’s from the sale of the debt. (Enter your answer in millions of dollars and round to 2 decimal places.)

Net Proceeds to Renee's $ _____________m

How many bonds will Renee’s Boutique need to sell in order to receive the $58.20 million it needs? (Do not round intermediate calculations and round your final answer to the nearest whole number.)

Number of Bonds = 63261
Answer #1

a.) Net Proceeds from the sale of Debt = $58.2 mn

Net Proceeds = Total Proceeds - Underwriters Commission.

Net Proceeds are required fund for Expansion.

b.) Total Proceeds = Required Fund / (1 - underwriter's spread) = 58,200,000 / (1 - .08) = $63,260,869.57 or $63.26 mn

Number of Bonds = Total Proceeds / Offer Price = 63,260,869.57 / 1000 = 63260.8 or 63261